# Analog In-memory Training on General Non-ideal Resistive Elements: The Impact of Response Functions

### v4

@@ Abstract
As the economic and environmental costs of training and deploying large vision or language models increase dramatically, analog in-memory computing (AIMC) emerges as a promising energy-efficient solution. However, the training perspective, especially its training dynamics, is underexplored. In AIMC hardware, the trainable weights are represented by the conductance of resistive elements and updated using consecutive electrical pulses. While the conductance changes by a constant in response to each pulse, in reality, the change is scaled by asymmetric and non-linear response functions, leading to a non-ideal training dynamics. This paper provides a theoretical foundation for gradient-based training on AIMC hardware with non-ideal response functions. We demonstrate that asymmetric response functions negatively impact Analog SGD by imposing an implicit penalty on the objective. To overcome the issue, we propose residual learning algorithm, which provably converges exactly to a critical point by solving a bilevel optimization problem. We demonstrate that the proposed method can be extended to address other hardware imperfections, such as limited response granularity. As we know, it is the first paper to investigate the impact of a class of generic non-ideal response functions. The conclusion is supported by simulations validating our theoretical insights.
@@ 2 Analog Training with Generic Response Functions
Gradient-based training algorithms on AIMC hardware. In (5), the desired update $\Delta W_{k}$ varies based on different algorithms. Replacing $\Delta W_{k}$ with the stochastic gradient $\nabla f(W_{k};\xi_{k})$, we obtain the dynamics of Analog SGD shown in (2). This update is reduced to the mathematical form for linear response functions in [21] as a special case; see Appendix B for details. Response function class. Before proceeding to the study of response functions, we first define the response function class that interests us. Since the behavior of resistive elements is always governed by physical laws, the function class should reflect certain crucial physical properties.
We first introduce a commonly observed saturation property across a wide range of resistive elements. Resistive elements get saturated when they keep receiving the same pulses to avoid reaching arbitrarily high or low conductance states. Constrained by that, the conductance of the resistive element with $q_{+}(\,\cdot\,)$ and $q_{-}(\,\cdot\,)$ is bounded inside a dynamic range $\tau^{\min},\tau^{\max}]$ where $\tau^{\min}$ and $\tau^{\max}$ are the saturation points with zero responses, i.e. $q_{+}(\tau^{\max})=q_{-}(\tau^{\min})=0$. Near the saturation points, the asymmetric issue is significant, so the update in one direction is suppressed, significantly impacting convergence. On the contrary, if a point $w^{\diamond}$ satisfies $q_{-}(w^{\diamond})=q_{+}(w^{\diamond})$, the analog update behaves like a digital update. We refer to $w^{\diamond}$ as symmetric point. Symmetric points are typically located in the interior of the dynamic range and are far from saturation.
Stacking all $w^{\diamond}$ into a vector $W^{\diamond}\in{\mathbb{R}}^{D}$. Observe that the function $G(W)$ is large near the saturation points while almost zero around $W^{\diamond}$, implying it can reflect the saturation degree. At the same time, $F(W)$ is the average of the response functions in two directions. As we will see in Sections 3.2 and 4, the ratio $\frac{G(W)}{\sqrt{F(W)}}$ plays a critical role in the convergence behaviors.
@@
Besides saturation, the function class should also enjoy other properties. First, the conductance increases when receiving a positive pulse and vice versa, leading to positive response functions. On top of that, we also assume the response functions are differentiable (and hence continuous) for mathematical tractability. Taking all into account, we define the following response function class.
@@
$q_{+}(\,\cdot\,)$ and $q_{-}(\,\cdot\,)$ with dynamic range $\tau^{\min},\tau^{\max}]$ satisfy
  • (Positive-definiteness) $q_{+}(w)>0,\forall w<\tau^{\max}$ and $q_{-}(w)>0,\forall w>\tau^{\min}$;
  • (Differentiable) $q_{+}(\,\cdot\,)$ and $q_{-}(\,\cdot\,)$ are differentiable;
  • (Saturation) $q_{+}(\tau^{\max})=q_{-}(\tau^{\min})=0$.
@@ 3.1 Implicit penalty
Assume ${\mathbb{E}}_{\xi_{k}}[|\nabla f(W_{k};\xi_{k})-{\mathbb{E}}_{\xi}[\nabla f(W_{k};\xi)]|]$ is a constant $\Sigma$ and suppose that $D=1$. Analog SGD implicitly optimizes the following penalized objective
@@ 3.2 Inexact Convergence of Analog SGD under generic devices
The stochastic gradient is unbiased and has bounded variance $\sigma^{2}$.
Assumption 12 are standard in non-convex optimization [38, 21]. Additionally, similar to the setting in [21], we also assume that the saturation degree is bounded, i.e., all response functions are positive.
Assumption 3 (Bounded saturation).
There exists a constant $R_{\min}>0$ such that $\min\{Q_{+}(W)\odot Q_{-}(W)\}\!>\!R_{\min}$.
Assumption 3 requires that $W_{k}$ remains far from saturation points, which is a mild assumption in actual training. This paper considers the average squared norm of the gradient as the convergence metric, given by $E^{\texttt{ASGD}}_{K}:=\frac{1}{K}\sum_{k=0}^{K-1}\left\|\nabla f(W_{k})\right\|^{2}$. Now, we establish the convergence of Analog SGD.
@@ 4 Mitigating Implicit Penalty by Residual Learning
Under Assumptions 13 and 4, with the learning rate $\alpha=O\left(\sqrt{{1}/{\sigma^{2}K}}\right)$, $\beta=O(\alpha\gamma^{3/2})$, it holds for Residual Learning that
@@
Under Assumption 5 and the conditions in Theorem 3, if $\gamma\geq\Omega(R_{\min}^{-1/5})$, it holds that $E^{\texttt{RL}}_{K}\leq\ O\left(\sqrt{{\sigma^{2}L}/{K}}\right)$.
@@ Appendix A Literature Review
Tiki-Taka  [Corollary 1] $\neq 0$ $O\left(\sqrt{\frac{1}{K}}\frac{1}{R^{\texttt{RL}}_{\min}}\right)$ $O\left(\sqrt{\frac{1}{K}}\frac{1}{1-P_{\max}^{2}/\tau^{2}}\right)$
Table 2: Comparison between our paper and [21]. Mixing-coefficient $\gamma$ is a hyper-parameter of Tiki-Taka. “Generic response” and “Linear response” columns are the convergence rates in the corresponding settings. $K$ represents the number of iterations. $R^{\texttt{RL}}_{\min}$ and $P_{\max}^{2}/\tau^{2}<1$ measure the saturation while the former one reduces to the latter on linear response functions.
@@ Appendix B Relation with the result in [21]
$\min\{R(W_{k})\}=$ $\ \min\{Q_{+}(W_{k})\odot Q_{-}(W_{k})\}=1-\left(\frac{\|W_{k}\|_{\infty}}{\tau}\right)^{2}$ (17)
$R^{\texttt{ASGD}}_{\min}=$ $\ \min\{R(W_{k})\}=1-\left(\frac{W_{\max}}{\tau}\right)^{2}.$ (18)
@@
With the linear response, if we also assume the bounded saturation of $P_{k}$ by letting $\|P_{k}\|_{\infty}\leq P_{\max}$, the minimal average response function is given by $R^{\texttt{RL}}_{\min}=1-\left(\frac{P_{\max}}{\tau}\right)^{2}$. The upper bound in Corollary 1 becomes
@@ Appendix G Proof of Theorem 1: Implicit Bias of Analog Training
In this section, we provide a full version of Theorem 1. Before that, we formally define the accumulated asymmetric function $R_{c}(W):{\mathbb{R}}^{D}\to{\mathbb{R}}^{D}$ element-wise. Let $R(W):=\frac{G(W)}{F(W)}$ be the asymmetric ratio and $R_{c}(W)$ is defined by
$[R_{c}(W)]_{d}:=\int_{\tau_{i}^{\min}}^{[W]_{d}}~[R(W)]_{d}~{\mathrm{d}[W]_{d}}$ (40)
which satisfies $\nabla\left\langle\Sigma,R_{c}(W)\right\rangle=\Sigma\odot R(W)$ for any $\Sigma\in{\mathbb{R}}^{D}$. Since we do not further assume stronger properties for response functions, like monotonicity, it is hard to provide strong claims on the shape of $R(W)$ or $R_{c}(W)$. Here we provide the expression of $R_{c}(W)$ for the linear response functions $Q_{+}(W)=1-\frac{W}{\tau}$, $Q_{-}(W)=1+\frac{W}{\tau}$. In this case, $F(W)=1$ and $G(W)=\frac{W}{\tau}$ based on definition (4); and hence $R(W)=\frac{G(W)}{F(W)}=\frac{W}{\tau}$. Accordingly, the accumulated asymmetric function is given by
@@
which is a weighted $\ell_{2}$ norm regularization term. In the scalar case, it reduces to
$\min_{W}~f_{\Sigma}(W):=f(W)+\frac{\Sigma}{2\tau}\|W\|^{2}.$ (43)
If the ratio $R(W)$ is monotonic at each coordinate, $R_{c}(W)$ reaches its minimum at $W^{\diamond}$. Therefore, $R_{c}(W)$ has no impact on the convergence only when the optimal point of $f(W)$ is $W^{\diamond}$.
Theorem 5 (Implicit Penalty, full version of Theorem 1).
Let $T(w)$ denote the effective update of Analog SGD.
$T(W):=\left|\frac{{\mathbb{E}}_{\xi}\left[U_{q}(W,-\alpha f^{\prime}(W;\xi))\right]-W}{\alpha}\right|=\left|{\mathbb{E}}_{\xi}[f^{\prime}(w;\xi)]\odot F(W)-{\mathbb{E}}_{\xi}[|f^{\prime}(W;\xi)|]\odot G(W)\right|.$ (44)
Analog SGD implicitly optimizes the following penalized objective
$\min_{W}~f_{\Sigma}(W):=f(W)+\left\langle\Sigma,R_{c}(W)\right\rangle$ (45)
in the sense that there exists a point $W^{S}$ given by
$W^{S}:=(\nabla^{2}f(W^{*})-\nabla R(W^{\diamond})\Sigma)^{-1}{(\nabla^{2}f(W^{*})~W^{*}-\nabla R(W^{\diamond})\Sigma~W^{\diamond})}{}$ (46)
such that $\|\nabla f_{\Sigma}(W^{S})\|\leq O((W^{\diamond}-W^{*})^{2})$ and $T(W^{S})\leq O((W^{\diamond}-W^{*})^{2})$. Both $T(W^{S})$ and $\|\nabla f_{\Sigma}(W^{S})\|$ are significantly smaller than $T(W^{\diamond})=O(|W^{\diamond}-W^{*}|)$ and $T(W^{*})=O(|W^{\diamond}-W^{*}|)$ when $W^{\diamond}$ is close to $W^{*}$.
If $W$ is a scalar, i.e. $D=1$, (46) reduces to (43)
$\min_{W}~f_{\Sigma}(W):=f(W)+\frac{\Sigma}{2\tau}\|W\|^{2}$ (47)
with its solution

### v5

@@ Abstract
As the economic and environmental costs of training and deploying large vision or language models increase dramatically, analog in-memory computing (AIMC) emerges as a promising energy-efficient solution. However, the training perspective, especially its training dynamics, is underexplored. In AIMC hardware, the trainable weights are represented by the conductance of resistive elements and updated using consecutive electrical pulses. While the conductance changes by a constant in response to each pulse, in reality, the change is scaled by asymmetric and non-linear response functions, leading to a non-ideal training dynamics. This paper provides a theoretical foundation for gradient-based training on AIMC hardware with non-ideal response functions. We demonstrate that asymmetric response functions negatively impact Analog SGD by imposing an implicit penalty on the objective. To address the issue, we propose residual learning algorithm, which provably converges exactly to a critical point by solving a bilevel optimization problem. We demonstrate that the proposed method can be extended to address other hardware imperfections, such as limited response granularity. As we know, it is the first paper to investigate the impact of a class of generic non-ideal response functions. The conclusion is supported by simulations validating our theoretical insights.
@@ 2 Analog Training with Generic Response Functions
Gradient-based training algorithms on AIMC hardware. In (5), the desired update $\Delta W_{k}$ varies based on different algorithms. Replacing $\Delta W_{k}$ with the stochastic gradient $\nabla f(W_{k};\xi_{k})$, we obtain the dynamics of Analog SGD shown in (2). This update is reduced to the mathematical form for linear response functions in [21] as a special case; see Appendix B for details.
Response function class. Before proceeding to the study of response functions, we first define the response function class that interests us. Since the behavior of resistive elements is always governed by physical laws, the function class should reflect certain crucial physical properties.
The most crucial property of the response functions is the asymmetric update, i.e., $q_{-}(w)\neq q_{+}(w)$ for most of $w$. Specifically, if a point $w^{\diamond}$ satisfies $q_{-}(w^{\diamond})=q_{+}(w^{\diamond})$, we say $w^{\diamond}$ is a symmetric point. Stacking all $w^{\diamond}$ into a vector $W^{\diamond}\in{\mathbb{R}}^{D}$. Observe that the function $G(W)$ is large if $q_{-}(w)$ and $q_{+}(w)$ are significantly different, while it is almost zero around $W^{\diamond}$. At the same time, $F(W)$ is the average of the response functions in two directions. As we will see in Sections 3.2 and 4, the ratio $\frac{G(W)}{\sqrt{F(W)}}$ plays a critical role in the convergence behaviors.
@@
In addition to the asymmetric update, the function class should possess other properties. First, the conductance increases upon receipt of a positive pulse, and vice versa, resulting in positive response functions. In addition, we assume that the response functions are differentiable (and hence continuous) for mathematical tractability. Taking all factors into account, we define the following class of response functions.
@@
$q_{+}(\,\cdot\,)$ and $q_{-}(\,\cdot\,)$ satisfy
  • (Positive-definiteness) There exist positive constants $q_{\min}>0$ and $q_{\max}>0$ such that $q_{\min}\leq q_{+}(w)\leq q_{\max}$ and $q_{\min}\leq q_{-}(w)\leq q_{\max},\forall w$; and,
  • (Differentiable) The response functions $q_{+}(\,\cdot\,)$ and $q_{-}(\,\cdot\,)$ are differentiable.
@@ 3.1 Implicit penalty
Suppose $W^{*}$ is the unique minimizer of problem (1). Let $\Sigma:={\mathbb{E}}_{\xi}[|\nabla f(W^{*};\xi)|]\in{\mathbb{R}}^{D}$. Analog SGD implicitly optimizes the following penalized objective
@@ 3.2 Inexact Convergence of Analog SGD under generic devices
The stochastic gradient is unbiased and has bounded variance $\sigma^{2}$. i.e., ${\mathbb{E}}_{\xi}[\nabla f(W;\xi)]=\nabla f(W)$ and ${\mathbb{E}}_{\xi}[\|\nabla f(W;\xi)-\nabla f(W)\|^{2}]\leq\sigma^{2}$.
Assumption 12 are standard in non-convex optimization [38]. This paper considers the average squared norm of the gradient as the convergence metric, given by $E^{\texttt{ASGD}}_{K}:=\frac{1}{K}\sum_{k=0}^{K-1}\left\|\nabla f(W_{k})\right\|^{2}$. Now, we establish the convergence of Analog SGD.
@@ 4 Mitigating Implicit Penalty by Residual Learning
Under Assumptions 13, with the learning rate $\alpha=O\left(\sqrt{{1}/{\sigma^{2}K}}\right)$, $\beta=O(\alpha\gamma^{3/2})$, it holds for Residual Learning that
@@
Under Assumption 4 and the conditions in Theorem 3, if $\gamma\geq\Omega(q_{\min}^{-2/5})$, it holds that $E^{\texttt{RL}}_{K}\leq\ O\left(\sqrt{{\sigma^{2}L}/{K}}\right)$.
@@ Appendix A Literature Review
Tiki-Taka  [Corollary 1] $\neq 0$ $O\left(\sqrt{\frac{1}{K}}\frac{1}{M^{\texttt{RL}}_{\min}}\right)$ $O\left(\sqrt{\frac{1}{K}}\frac{1}{1-P_{\max}^{2}/\tau^{2}}\right)$
Table 2: Comparison between our paper and [21]. Mixing-coefficient $\gamma$ is a hyper-parameter of Tiki-Taka. “Generic response” and “Linear response” columns are the convergence rates in the corresponding settings. $K$ represents the number of iterations. $M^{\texttt{RL}}_{\min}$ and $P_{\max}^{2}/\tau^{2}<1$ measure the saturation while the former one reduces to the latter on linear response functions.
@@ Appendix B Relation with the result in [21]
$q_{\max}=$ $\ 1+\frac{W_{\max}}{\tau},\quad q_{\min}=1-\frac{W_{\max}}{\tau},$ (17)
$\min\{M(W_{k})\}=$ $\ \min\{Q_{+}(W_{k})\odot Q_{-}(W_{k})\}=1-\left(\frac{\|W_{k}\|_{\infty}}{\tau}\right)^{2}$ (18)
$M^{\texttt{ASGD}}_{\min}=$ $\ \min\{M(W_{k})\}=1-\left(\frac{W_{\max}}{\tau}\right)^{2}.$ (19)
@@
With the linear response, if we also assume the bounded saturation of $P_{k}$ by letting $\|P_{k}\|_{\infty}\leq P_{\max}$, the minimal average response function is given by $M^{\texttt{RL}}_{\min}=1-\left(\frac{P_{\max}}{\tau}\right)^{2}$. The upper bound in Corollary 1 becomes
@@ Appendix G Proof of Theorem 1: Implicit Bias of Analog Training
In this section, we provide a full version of Theorem 1. Before that, we introduce the asymmetry ratio and its element-wise anti-derivative. Since $F(\cdot)$ and $G(\cdot)$ act element-wise, we define $R(\,\cdot\,):{\mathbb{R}}^{D}\to{\mathbb{R}}^{D}$ and $R_{c}(\,\cdot\,):{\mathbb{R}}^{D}\to{\mathbb{R}}^{D}$ element-wise by
$R(W):=\frac{G(W)}{F(W)}\quad\text{and}\quad[R_{c}(W)]_{d}:=\int_{[W^{\diamond}]_{d}}^{[W]_{d}}[R(W^{\prime})]_{d}\,{\mathrm{d}[W^{\prime}]_{d}},\quad d\in[D],$ (41)
where the division is taken element-wise. It holds that $\frac{{\mathrm{d}}}{{\mathrm{d}[W]_{d}}}[R_{c}(W)]_{d}=[R(W)]_{d}$ for each coordinate $d\in[D]$, and $\nabla\left\langle C,R_{c}(W)\right\rangle=C\odot R(W)$ for any vector $C\in{\mathbb{R}}^{D}$. Consequently, if $R(W)$ is strictly monotonic, $R_{c}(W)\geq 0$, and it reaches its minimum value at the symmetric point $W^{\diamond}$ where $R(W^{\diamond})=0$.
Define the scaled effective update of Analog SGD at $W$ as
$T(W):={\mathbb{E}}_{\xi}[\nabla f(W;\xi)]+{\mathbb{E}}_{\xi}[|\nabla f(W;\xi)|]\odot\frac{G(W)}{F(W)}.$ (42)
A direct verification shows that if $T(W_{k})=0$, then Analog SGD (2) (equivalently, (5) with $\Delta W_{k}=-\alpha\nabla f(W_{k};\xi_{k})$) is stable at $W_{k}$ in conditional expectation, i.e., ${\mathbb{E}}_{\xi_{k}}[W_{k+1}]=W_{k}$.
Rather than characterizing the exact limit of Analog SGD, we establish a local statement showing the existence of a point that is simultaneously (i) nearly critical for $f_{\Sigma}$; and, (ii) nearly stationary for the mean Analog SGD dynamics.
Theorem 5 (Implicit Penalty, full version of Theorem 1).
Suppose $W^{*}$ is the unique minimizer of problem (1). Define
${\tilde{W}}^{*}:=\Big(\nabla^{2}f(W^{*})+\text{Diag}(\Sigma)\,\nabla R(W^{\diamond})\Big)^{-1}{\Big(\nabla^{2}f(W^{*})~W^{*}+\text{Diag}(\Sigma)\,\nabla R(W^{\diamond})W^{\diamond}\Big)}{}$ (43)
where $\text{Diag}(M)\in\mathbb{R}^{D\times D}$ denotes the diagonal matrix whose diagonal entries are given by the vector $M\in\mathbb{R}^{D}$. Let $\Sigma:={\mathbb{E}}_{\xi}[|\nabla f(W^{*};\xi)|]\in{\mathbb{R}}^{D}$ be the element-wise first moment of stochastic gradients at $W^{*}$. Analog SGD implicitly optimizes
$\min_{W\in{\mathbb{R}}^{D}}~f_{\Sigma}(W):=f(W)+\left\langle\Sigma,R_{c}(W)\right\rangle$ (44)
in the sense that as $\|W^{\diamond}-W^{*}\|\to 0$,
$\frac{\|\nabla f_{\Sigma}({\tilde{W}}^{*})\|}{\min\{\|\nabla f_{\Sigma}(W^{*})\|,\|\nabla f_{\Sigma}(W^{\diamond})\|\}}\to 0\quad\text{and}\quad\frac{\|T({\tilde{W}}^{*})\|}{\min\{\|T(W^{*})\|,\|T(W^{\diamond})\|\}}\to 0.$ (45)
Proof of Theorem 5.
By the definition of $\tilde{W}$$^{*}$, it holds that
$\|W^{\diamond}-{\tilde{W}}^{*}\|$= $\ \|(\nabla^{2}f(W^{*})+\text{Diag}(\Sigma)\,\nabla R(W^{\diamond}))^{-1}\nabla^{2}f(W^{*})~(W^{*}-W^{\diamond})\|$
$\ =\Theta(\|W^{\diamond}-W^{*}\|)$ (46)
$\|W^{*}-{\tilde{W}}^{*}\|$= $\ \|(\nabla^{2}f(W^{*})+\text{Diag}(\Sigma)\,\nabla R(W^{\diamond}))^{-1}\text{Diag}(\Sigma)\,\nabla R(W^{\diamond})(W^{*}-W^{\diamond})\|$
$\ =\Theta(\|W^{\diamond}-W^{*}\|).$ (47)
We separately show the two parts of Theorem 5. We first show that $\|T({\tilde{W}}^{*})\|\leq{\mathcal{O}}(\|W^{\diamond}-W^{*}\|^{2})$, and $\|T(W^{\diamond})\|=\Theta(\|W^{\diamond}-W^{*}\|)$ and $\|T(W^{*})\|=\Theta(\|W^{\diamond}-W^{*}\|)$. It reaches the first limit in (45). Similarly, we show the second limit by showing that $\|\nabla f_{\Sigma}({\tilde{W}}^{*})\|\leq{\mathcal{O}}(\|W^{\diamond}-W^{*}\|^{2})$, and $\|\nabla f_{\Sigma}(W^{\diamond})\|=\Theta(\|W^{\diamond}-W^{*}\|)$ and $\|\nabla f_{\Sigma}(W^{*})\|=\Theta(\|W^{\diamond}-W^{*}\|)$.
(Step 1a) Proof of $\|\nabla f_{\Sigma}({\tilde{W}}^{*})\|\leq{\mathcal{O}}(\|W^{\diamond}-W^{*}\|^{2})$. The gradient of $f_{\Sigma}(W)$ is given by
$\nabla f_{\Sigma}(W)=\nabla f(W)+\Sigma\odot R(W).$ (48)
[…]
$\frac{G({\tilde{W}}^{*})}{F({\tilde{W}}^{*})}$= $\ \nabla R(W^{\diamond})({\tilde{W}}^{*}-W^{\diamond})+{\mathcal{O}}(({\tilde{W}}^{*}-W^{\diamond})^{2}),$ (50)
where ${\mathcal{O}}(({\tilde{W}}^{*}-W^{*})^{2})$ and ${\mathcal{O}}(({\tilde{W}}^{*}-W^{\diamond})^{2})$ are vectors with norms $\|{\tilde{W}}^{*}-W^{*}\|^{2}$ and $\|{\tilde{W}}^{*}-W^{\diamond}\|^{2}$, respectively. We bound the gradient of the penalized objective as follows
$\ \|\nabla f_{\Sigma}({\tilde{W}}^{*})\|=\left\|\nabla f({\tilde{W}}^{*})+\Sigma\odot\frac{G({\tilde{W}}^{*})}{F({\tilde{W}}^{*})}\right\|$ (51)
$=$ $\ \left\|\nabla^{2}f(W^{*})({\tilde{W}}^{*}-W^{*})+{\mathcal{O}}(({\tilde{W}}^{*}-W^{*})^{2})+\Sigma\odot(\nabla R(W^{\diamond})({\tilde{W}}^{*}-W^{\diamond}))+{\mathcal{O}}(({\tilde{W}}^{*}-W^{\diamond})^{2})\right\|$
$=$ $\ \|{\mathcal{O}}(({\tilde{W}}^{*}-W^{*})^{2})+{\mathcal{O}}(({\tilde{W}}^{*}-W^{\diamond})^{2})\|={\mathcal{O}}(\|W^{*}-W^{\diamond}\|^{2})$
where the last inequality holds by (G) and (G).
(Step 1b) Proof of $\|T(W^{\diamond})\|=\Theta(\|W^{\diamond}-W^{*}\|)$ and $\|T(W^{*})\|=\Theta(\|W^{\diamond}-W^{*}\|)$.
$\|\nabla f_{\Sigma}(W^{\diamond})\|=$ $\ \left\|\nabla f(W^{\diamond})+\Sigma\odot R(W^{\diamond})\right\|=\left\|\Sigma\odot R(W^{*})\right\|=\Theta(\|W^{*}-W^{\diamond}\|),$ (52)
$\|\nabla f_{\Sigma}(W^{*})\|=$ $\ \left\|\nabla f(W^{*})+\Sigma\odot R(W^{*})\right\|=\left\|\nabla f(W^{\diamond})\right\|=\Theta(\|W^{*}-W^{\diamond}\|).$ (53)
[…]
$\ \|T({\tilde{W}}^{*})\|=\left\|{\mathbb{E}}_{\xi}[\nabla f({\tilde{W}}^{*};\xi)]-{\mathbb{E}}_{\xi}[|\nabla f({\tilde{W}}^{*};\xi)|]\odot\frac{G({\tilde{W}}^{*})}{F({\tilde{W}}^{*})}\right\|$ (54)
$\leq$ $\ \left\|\nabla f({\tilde{W}}^{*})-{\mathbb{E}}_{\xi}[|\nabla f({\tilde{W}}^{*};\xi)|]\odot\frac{G({\tilde{W}}^{*})}{F({\tilde{W}}^{*})}\right\|$
$\leq$ $\ \left\|\nabla f({\tilde{W}}^{*})-{\mathbb{E}}_{\xi}[|\nabla f(W^{*};\xi)|]\odot\frac{G({\tilde{W}}^{*})}{F({\tilde{W}}^{*})}\right\|+\left\|({\mathbb{E}}_{\xi}[|\nabla f(W^{*};\xi)|]-{\mathbb{E}}_{\xi}[|\nabla f({\tilde{W}}^{*};\xi)|])\odot\frac{G({\tilde{W}}^{*})}{F({\tilde{W}}^{*})}\right\|$
$=$ $\ \|\nabla f_{\Sigma}({\tilde{W}}^{*})\|+\left\|({\mathbb{E}}_{\xi}[|\nabla f(W^{*};\xi)|]-{\mathbb{E}}_{\xi}[|\nabla f({\tilde{W}}^{*};\xi)|])\odot\frac{G({\tilde{W}}^{*})}{F({\tilde{W}}^{*})}\right\|$
The first term in the RHS (RHS) of (54) is bounded by (51). By applying $||x|-|y||\leq|x-y|$ for any $x,y\in{\mathbb{R}}$ at all components, the second term in the RHS of (54) is bounded by
$\ \left\|({\mathbb{E}}_{\xi}[|\nabla f(W^{*};\xi)|]-{\mathbb{E}}_{\xi}[|\nabla f({\tilde{W}}^{*};\xi)|])\odot\frac{G({\tilde{W}}^{*})}{F({\tilde{W}}^{*})}\right\|$ (55)
$\leq$ $\ \left\|{\mathbb{E}}_{\xi}[|\nabla f(W^{*};\xi)-\nabla f({\tilde{W}}^{*};\xi)|]\odot\frac{G({\tilde{W}}^{*})}{F({\tilde{W}}^{*})}-\frac{G(W^{\diamond})}{F(W^{\diamond})}\right\|$
$\leq$ $\ \left\|{\mathbb{E}}_{\xi}[|\nabla f(W^{*};\xi)-\nabla f({\tilde{W}}^{*};\xi)|]\right\|\left\|\frac{G({\tilde{W}}^{*})}{F({\tilde{W}}^{*})}-\frac{G(W^{\diamond})}{F(W^{\diamond})}\right\|$
$\leq$ $\ {\mathbb{E}}_{\xi}\left[\left\|\nabla f(W^{*};\xi)-\nabla f({\tilde{W}}^{*};\xi)\right\|\right]\left\|\frac{G({\tilde{W}}^{*})}{F({\tilde{W}}^{*})}-\frac{G(W^{\diamond})}{F(W^{\diamond})}\right\|$
$\leq$ $\ {\mathcal{O}}(\|{\tilde{W}}^{*}-W^{*}\|\|{\tilde{W}}^{*}-W^{\diamond}\|)={\mathcal{O}}(\|W^{*}-W^{\diamond}\|^{2}).$
Plugging back (51) and (55) into (54) shows $T({\tilde{W}}^{*})\leq{\mathcal{O}}((W^{\diamond}-W^{*})^{2})$.
(Step 2b) Proof of $\|T(W^{\diamond})\|=\Theta(\|W^{\diamond}-W^{*}\|)$ and $\|T(W^{*})\|=\Theta(\|W^{\diamond}-W^{*}\|)$.
$\|T(W^{\diamond})\|=$ $\ \left\|{\mathbb{E}}_{\xi}[\nabla f(W^{\diamond};\xi)]-{\mathbb{E}}_{\xi}[|\nabla f(W^{\diamond};\xi)|]\odot\frac{G(W^{\diamond})}{F(W^{\diamond})}\right\|=\left\|\nabla f(W^{\diamond})\right\|$ (56)
$=$ $\ \Theta(\|W^{*}-W^{\diamond}\|),$
$\|T(W^{*})\|=$ $\ \left\|{\mathbb{E}}_{\xi}[\nabla f(W^{*};\xi)]-{\mathbb{E}}_{\xi}[|\nabla f(W^{*};\xi)|]\odot\frac{G(W^{*})}{F(W^{*})}\right\|=\left\|\Sigma\odot R(W^{*})\right\|$ (57)
$=$ $\ \Theta(\|W^{*}-W^{\diamond}\|).$
Now we complete the proof. ∎
Special case with $D=1$.
Before proceeding to the next section, we also present a special case where the response functions are linear, i.e., $Q_{+}(W)=1-\frac{W}{\tau}$, $Q_{-}(W)=1+\frac{W}{\tau}$. $F(W)=1$ and $G(W)=\frac{W}{\tau}$ based on definition (4); and hence $R(W)=\frac{G(W)}{F(W)}=\frac{W}{\tau}$. Accordingly, the accumulated asymmetric function is given by
@@
which is a weighted $\ell_{2}$ norm regularization term. Furthermore, if $W$ is a scalar, i.e., $D=1$, (44) reduces to $\min_{W}~f_{\Sigma}(W):=f(W)+\frac{\Sigma}{2\tau}\|W\|^{2}$, which is a $\ell_{2}$-regularized problem with an approximated solution